Grade 12th passMechanics

In an isothermal expansion of work done by a system is -x is heat lost from the system or given to the system

Anoushka gowda
10 Years agoGrade 12th pass

1 Answer

ADITI SINGH
10 Years ago
Using the ideal gas law, a relationship between p and V is found to be p = nRT V . From the graph,it follows that the pressure is inversely proportional to the volume of the gas, e.g. p/p0 = 1/4 when V/V0 = 4. It can be seen from the equation above that this can only be true if the temperature of the gas remains constant. The curve OC is then representative of a isothermal process.